INFRARED DIODE LASER SPECTROSCOPY OF N_{2}O CLUSTERS IN THE N-O STRETCHING REGION

\underline{M. DEHGHANY}, M. AFSHARI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, 2500 University Drive NW, Calgary, ALberta T2N 1N4, Canada; A.R.W. MCKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A OR6, Canada.

The high resolution infrared spectrum of the polar N_{2}O dimer has been observed in the region of N-O stretching fundamental using a tunable diode laser to probe a pulsed supersonic slit jet. An a/b type hybrid band has been assigned and fitted to a planar asymmetric top with slipped parallel structure. The vibrational origin is measured to be 1290.2 cm^{−1} showing a blue shift of 5.3 cm^{−1} from the monomer band origin and 10.5 cm^{−1} from the band origin of the non-polar isomer. The band for the non-polar isomer has also been remeasured and analysed in improved detail. In addition, two trimer bands of N_{2}O, located at 1276.56 cm^{−1} and 1285.78 cm^{−1}, have been observed and assigned with a/c and a-type transitions, respectively.